OpenStudy (anonymous):

Help Please ? Can Some Check If I'm Right and Correct If I'm Wrong? Just 3 Questions :) Thank You In Advance

OpenStudy (anonymous):

I can try (:

OpenStudy (anonymous):

Appreciate It Very much, Ill Post Them :)

OpenStudy (anonymous):

What is the equation of the graph below? y = - (x - 3)2 + 1 y = - (x + 3)2 + 1 y = (x - 3)2 - 1 y = (x + 3)2 - 1 <---------------------- My answer

OpenStudy (anonymous):

Nm Its Only 2 Quetions I need Help With :)

OpenStudy (anonymous):

This one and another one

OpenStudy (anonymous):

You were right! It is D!

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

You're welcome, what is the other one?